# Robust Design and Evaluation of

### v5

@@ Abstract
Predictive algorithms inform consequential decisions in settings where the outcome is selectively observed given choices made by human decision makers. We propose a unified framework for the robust design and evaluation of predictive algorithms in selectively observed data. We impose general assumptions on how much the outcome may vary on average between unselected and selected units conditional on observed covariates and identified nuisance parameters, formalizing popular empirical strategies for imputing missing data such as proxy outcomes and instrumental variables. We develop debiased machine learning estimators for the bounds on a large class of predictive performance estimands, such as the conditional likelihood of the outcome, a predictive algorithm’s mean square error, true/false positive rate, and many others, under these assumptions. In an administrative dataset from a large Australian financial institution, we illustrate how varying assumptions on unobserved confounding leads to meaningful changes in default risk predictions and evaluations of credit scores across sensitive groups.
@@ 1 Introduction
Predictive algorithms inform high-stakes decisions by providing predictions about the likelihood of an outcome based on observed covariates. In designing and evaluating such tools, we often face a missing data or selective labels problem: the outcome is only selectively observed conditional on some choices made by human decision makers. For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released prior to their trial; but we only observe whether a past defendant failed to appear in court if a judge decided to release them (e.g., Kleinberg et al., 2018). Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether a past applicant defaulted if the financial institution approved them and they accepted the offered loan terms (e.g., Blattner and Nelson, 2021; Coston et al., 2021). [^1] This missing data problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., Mullainathan and Obermeyer, 2022), child protective services (e.g., Chouldechova et al., 2018), hiring (e.g., Li et al., 2020), education (Bergman et al., 2021), and tax audits (e.g., Black et al., 2022; Battaglini et al., 2022; Elzayn et al., 2023).
@@
Addressing this missing data problem is difficult as there may exist unobserved confounders that affect both the selection decision and the outcome. Judges’ pretrial release decisions and loan applicants’ decisions to accept offered loan terms likely depend on additional private information that is not fully captured by observed covariates. Naively ignoring such unobserved confounding would lead to inaccurate predictions and misleading evaluations of predictive algorithms. [^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston et al., 2020; Coston et al., 2021; Mishler et al., 2021; Mishler and Kennedy, 2021; Guerdan et al., 2023), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates. While this challenge is recognized, existing solutions are limited. Empirical researchers often rely on heuristics for imputing missing data based on observed outcomes among selected units, proxy outcomes (Chouldechova et al., 2018; Blattner and Nelson, 2021; Mullainathan and Obermeyer, 2022), or instrumental variable strategies tailored to particular evaluation estimands (Lakkaraju et al., 2017; Kleinberg et al., 2018; Arnold et al., 2020) and predictive algorithms that are linear in covariates (Arnold et al., 2024).
In this paper, we develop a framework for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding. We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters. Such restrictions formalize a wide variety of popular empirical strategies for addressing unobserved confounding and partially identify a large class of predictive estimands found in applied research studying predictive algorithms. We develop debiased machine learning estimators for the bounds on these quantities. Our framework produces informative bounds by incorporating observed covariates flexibly in our identifying assumptions and estimation procedures.
@@
We consider settings with data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from some joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision determined by past decision makers such as judges or loan applicants. The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$, where for example $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if they were released before trial or whether an applicant would default if they were granted a loan. [^3] We make the restriction that $Y_{i}^{*}$ is binary because the true outcome is binary in all empirical applications mentioned so far (see Section 2 for more details).
In this setting, empirical researchers typically have two objectives. First, they wish to construct predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$, and data-driven decision rules implied by the predictive algorithm. Second, when given an existing predictive algorithm, empirical researchers evaluate its overall performance
$\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}],$ (1)
as well as its positive class and negative class performance

### v6

@@ Abstract
Predictive algorithms inform consequential decisions in settings with selective labels: outcomes are observed only for units selected by past decision makers. This creates an identification problem under unobserved confounding — when selected and unselected units differ in unobserved ways that affect outcomes. We propose a framework for robust design and evaluation of predictive algorithms that bounds how much outcomes may differ between selected and unselected units with the same observed characteristics. These bounds formalize common empirical strategies including proxy outcomes and instrumental variables. Our estimators work across bounding strategies and performance measures such as conditional likelihoods, mean square error, and true/false positive rates. Using administrative data from a large Australian financial institution, we show that varying confounding assumptions substantially affects credit risk predictions and fairness evaluations across income groups.
@@ 1 Introduction
Predictive algorithms inform high-stakes decisions in pretrial release, consumer lending, health care, and many other domains. Designing and evaluating such tools is often complicated by the “selective labels” problem: outcomes are observed only for units selected by past decision makers. For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released before trial; but we only observe whether a defendant committed misconduct if a judge decided to release them (e.g., KLLLM(18); Rambachan(21); AngelovaEtAl(22)). Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether an applicant defaulted if the financial institution approved them and they accepted the loan terms (e.g., BlattnerNelson(21); CostonEtAl(21)-Rashomon). [^1] The selective labels problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., MullainathanObermeyer(21)), child protective services (e.g., ChouldechovaEtAl(18)), hiring (e.g., LiRaymondBergman(20)), education (BergmanEtAl(21)), and tax audits (e.g., BlackEtAl(22)-TaxAudits; BattagliniEtAl(22); ElzaynEtAl(23)).
@@
The selective labels problem creates an identification challenge under unobserved confounding–when selected and unselected units differ in ways not captured by observed covariates. Judges may release defendants based on courtroom interactions not in administrative data; loan applicants may accept offers based on competing credit options unavailable to researchers. Ignoring unobserved confounding leads to inaccurate predictions and misleading evaluations of predictive algorithms. [^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston(20)-Evaluation; CostonEtAl(21)-Rashomon; MishlerEtAl(21); mishler2021fade; guerdan2023counterfactual), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates. Existing solutions are limited. Applied researchers rely on informal heuristics—imputing missing outcomes using proxy variables (ChouldechovaEtAl(18); BlattnerNelson(21); MullainathanObermeyer(21)) or applying ad-hoc adjustments to observed outcomes among selected units. Recent work uses instrumental variables but typically focuses on specific estimands (lakkaraju2017selective; ArnoldDobbieHull(20)-discrimination_alg) or predictive algorithms that are linear in covariates (ArnoldEtAl(24)).
We develop a framework for robust design and evaluation of predictive algorithms under unobserved confounding. We bound how much outcomes may differ between selected and unselected units with the same observed characteristics. These bounds formalize common empirical strategies — including proxy outcomes and instrumental variables — and partially identify key predictive estimands such as conditional likelihoods, mean square error, calibration, and true/false positive rates. Our estimation approach works across multiple bounding strategies and performance measures, flexibly incorporating machine learning methods.
@@
We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision by past decision makers (e.g., judges or loan applicants). The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$; for example, $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if released before trial or whether an applicant would default if granted a loan.
Researchers have two objectives. First, design predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$ and construct data-driven decision rules. Second, evaluate existing predictive algorithms through performance measures, such as overall performance given by $\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$, and class-specific performance given by $\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]$ and $\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0]$, where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions. Alternative choices recover popular diagnostics such as mean square error, true positive rate, and false positive rate. Comparing these measures across sensitive attributes evaluates fairness properties (e.g., MitchellEtAl(19); barocas-hardt-narayanan).
Rather than assuming unconfounded selection, we bound the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using researcher-specified bounding functions that may depend on covariates and identified nuisance parameters. The confounding function measures how much unselected units differ from selected units with the same characteristics; bounding it formalizes assumptions about unobserved confounding severity without requiring a full selection model. Under these bounds, the conditional likelihood and predictive performance measures are partially identified.
Our framework encompasses popular strategies for addressing selective labels as special cases. When researchers observe proxy outcomes correlated with $Y_{i}^{*}$ — such as default on other credit products (BlattnerNelson(21)) or long-term health outcomes (MullainathanObermeyer(21); ChanGentzkowYu(21)) — this implies specific bounds. When researchers have instrumental variables like randomly assigned decision makers (lakkaraju2017selective; KLLLM(18)), this implies alternative bounds. Our framework accommodates other plausible identifying assumptions.
We apply this framework to two objectives. First, we develop estimators for the bounds on the conditional likelihood $P(Y_{i}^{*}=1\mid X_{i}=x)$ (e.g., the pretrial misconduct rate or default rate given covariates), enabling design of predictive algorithms robust to unobserved confounding. Building on advances in causal machine learning (NieWager(20); kennedy2020optimal), we propose pseudo-outcome regression estimators with sample splitting: estimate nuisance parameters on one fold, construct influence-function-based pseudo-outcomes, then regress them on covariates using machine learning on a second fold. A key innovation is an oracle inequality extending kennedy2020optimal’s pointwise analysis to integrated mean square error: our feasible estimators approximate an infeasible oracle up to a smoothed, doubly-robust product of first-stage errors. This applies to any regression method satisfying a mild stability condition, including random forests, kernel methods, and series estimators. We leverage this to characterize integrated mean square error convergence rates and establish regret bounds for plug-in decision rules, connecting estimation guarantees to decision-theoretic performance.
Second, we develop estimators for bounds on the performance of a given predictive algorithm $s(\cdot)$ under unobserved confounding. For overall performance measures $\mbox{perf}(s;\beta)$ — which nest mean square error, calibration, and precision — we construct debiased estimators using crossfitting and debiased machine learning (e.g., RobinsEtAl(08); ZhengvanderLaan(11); ChernozhukovEtAl(18)-DML). Under mild regularity conditions, these estimators are $\sqrt{n}$-consistent and jointly asymptotically normal. For class-specific performance $\mbox{perf}_{+}(s;\beta)$ and $\mbox{perf}_{-}(s;\beta)$ — such as true and false positive rates — we show sharp bounds correspond to optimal values of linear fractional programs with nuisances in both the objective and constraints. We estimate using two folds: estimate nuisances and bounding functions on the first fold, then solve the sample optimization program on the second fold. We establish partial double robustness: estimation error decomposes into sampling error, a doubly-robust product for objective nuisances, and constraint estimation errors entering through root mean square error. Proving this exploits the optimization structure of linear fractional programs. [^3] Since linear fractional programs arise in several related sensitivity analysis frameworks in causal inference (e.g., AronowLee(2013); MiratrixWagerZubizaretta(18); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21)), our results may be broadly applicable beyond the selective labels setting. Our estimation strategies accommodate multiple bounding strategies, flexibly incorporating domain-specific knowledge while maintaining statistical guarantees.
As an empirical illustration, we apply our framework to administrative data on personal loan applications from a large Australian bank. First, we construct confounding-robust credit risk scores under varying confounding assumptions. The resulting algorithms produce substantially different risk rankings that could meaningfully affect credit access. Second, inspired by recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate a benchmark credit score’s performance across income groups. An analysis ignoring unobserved confounding suggests the score is more accurate (in mean square error) for higher-income applicants. However, we show this disparity disappears at plausible levels of unobserved confounding, illustrating how researchers can assess the robustness of fairness conclusions.
Our framework prioritizes breadth and applicability. Rather than bespoke methods for each combination of bounding strategy and performance measure, we provide a unified approach for a variety of assumptions and performance measures. We deliver practical estimation procedures that enable researchers to robustly design and evaluate predictive algorithms under unobserved confounding.
Related Work:
Our analysis builds on foundational work on partial identification with selective labels (e.g., Manski(89); Manski(90); Manski(94)-SampleSelectionBounds). We generalize Manski(03)’s “approximate mean independence” assumption by incorporating covariates and nuisance parameters into the bounding functions, enabling flexible formalization of empirical heuristics using proxy outcomes and other strategies. See also diaz2013sensitivity; luedtke2015statistics; diaz2018sensitivity for similar bounding assumptions without covariates. We characterize identified sets for novel estimands arising in predictive algorithm evaluation.
Instrumental variable strategies are widely used to address unobserved confounding in this setting. lakkaraju2017selective and KLLLM(18) evaluate specific performance measures by returning a single point in the identified set. ArnoldDobbieHull(20)-discrimination_alg; ArnoldEtAl(24) extrapolate across instrument values using identification-at-infinity arguments. We instead derive covariate-dependent bounds following Manski(94)-SampleSelectionBounds, accommodating IVs as a special case while also encompassing other identifying strategies. [^4] In related work, LevisEtAl(23)-CovariateAssistedIV analyze covariate-assisted Balke-Pearl bounds on the average treatment effect. semenova2023generalized generalizes the Lee(09) bounds on the average treatment effect on the always takers, which requires a binary instrument that monotonically affects selection, to incorporate covariate information. Monotonicity is particularly unlikely to hold when the instrument arises from the random assignment of decision makers. See, for example, FrandsenLefgrenLeslie(19); ArnoldDobbieHull(20)-discrimination_bail; ChanGentzkowYu(21).
Our work relates to sensitivity analysis in causal inference, where researchers bound treatment effects under relaxations of unconfoundedness. Influential frameworks include the marginal sensitivity model (e.g., Tan(06); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); DornGuoKallus(21); KallusZhou(21); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); yadlowsky2018bounds), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)). These models bound how unobservables affect selection propensities—assumptions that may be difficult to formulate when decisions involve human judgment. We instead bound outcome differences, which may be more intuitive in applied settings. Nonetheless, these models imply specific bounding functions in our framework, as we discuss in Remark 2. chernozhukov2022long develop omitted variable bounds for linear functionals; this covers our overall performance measures but not class-specific performance.
2 Setup and Identification Framework
We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ are covariates and $D_{i}\in\{0,1\}$ is the decision. There is a “selective labels” problem: the observed outcome satisfies $Y_{i}=D_{i}\,Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$ (e.g., KLLLM(18)). We assume $0<\mathbb{P}(Y_{i}^{*}=1)<1$ so that we face a non-trivial prediction problem and that the decision satisfies strict overlap (i.e., there exists $\epsilon>0$ such that $\mathbb{P}(D_{i}=1\mid X_{i}=x)\geq\epsilon$ with probability one).
Example: lending decisions
We observe data on loan applications, where $X_{i}$ contains information reported on the application such as income and prior credit history, $D_{i}$ is whether the application was granted a loan, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the applicant defaulted if granted the loan. We study credit scores that predict whether a new applicant would default $Y_{i}^{*}=1$ (e.g., BlattnerNelson(21); DiMaggioEtAl(22)-Fintech; FusterEtAl(22)predictably). $\blacktriangle$
Example: pretrial release
We observe data on defendants, where $X_{i}$ contains information about the defendant such as their current charge and prior conviction history, $D_{i}$ is whether the defendant was released before trial, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the defendant committed pretrial misconduct if released. We study pretrial risk tools that predict whether a new defendant would commit pretrial misconduct $Y_{i}^{*}=1$ (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_alg; AngelovaEtAl(22)). $\blacktriangle$
Example: medical testing
We observe data on patients, where $X_{i}$ contains patient information such as vital signs at admission and prior medical history, $D_{i}$ is whether a doctor decided to conduct a diagnostic test such as a cardiac stress test or lab test, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the patient was diagnosed with a heart attack or bacterial infection. We study medical risk tools that predict whether a new patient suffered a heart attack or bacterial infection $Y_{i}^{*}=1$ (e.g., MullainathanObermeyer(21); HuangEtAl(22)). $\blacktriangle$
Let $\pi_{d}(x):=\mathbb{P}(D_{i}=d\mid X_{i}=x)$ for $d\in\{0,1\}$ be the propensity score and $\mu_{1}(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$. Averages of a random variable $V_{i}$ are $\mathbb{E}_{n}[V_{i}]:=n^{-1}\sum_{i=1}^{n}V_{i}$. We let $\|\cdot\|$ denote the appropriate $L_{2}$-norm by context, where $\|f\|=\left(\int f(v)^{2}d\mathbb{P}(v)\right)^{1/2}$ for a measurable function $f(\cdot)$ and $\|v\|=\left(\sum_{j=1}^{k}v_{j}^{2}\right)^{1/2}$ for a vector $v\in\mathbb{R}^{k}$.
2.1 Predictive Algorithms and Performance Measures
A predictive algorithm $s(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ predicts the probability $Y_{i}^{*}=1$ given covariates $X_{i}$, therefore estimating the conditional likelihood $\mu^{*}(x):=\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)$. Given a predictive algorithm, researchers evaluate its performance in two ways.
First, given a predictive algorithm, researchers construct decision rules $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ by applying a threshold rule; that is, defining $d(X_{i})=1\{s(X_{i})\geq\tau\}$ for some threshold $\tau\in[0,1]$. Alternative decision rules are evaluated based on their expected payoff
where each $u_{d,y^{*}}(\cdot)\geq 0$ for $d,y^{*}\in\{0,1\}^{2}$ specifies the known payoff associated with a combination of decision $D_{i}$ and outcome $Y_{i}^{*}$ at features $X_{i}$. Payoffs are normalized so that $\sum_{d,y^{*}\in\{0,1\}^{2}}u_{d,y^{*}}(x)=1$ for all $x\in\mathcal{X}$. In consumer lending, the profitability of approving customers that would not default $u_{1,0}(\cdot)$ may vary based on features like the requested loan size. In pretrial release, the societal benefits of releasing defendants that would not commit misconduct $u_{1,0}(\cdot)$ may vary based on the defendant’s age, charge severity, or prior history of pretrial misconduct (e.g., Baughman(17)).
Second, given a predictive algorithm $s(\cdot)$, researchers evaluate its performance over the full population or specific groups. The target parameters are its overall performance
and its positive class and negative class performance